# Pre-DPO: Improving Data Utilization in Direct Preference Optimization Using a Guiding Reference Model

### v2

@@ 4.2 Pre-DPO: improving data utilization in DPO using a guiding reference model
Notably, a model already undergone preference optimization contains information about the entire training process. More importantly, it reflects the outcomes that the initial policy can achieve through the available preference data. Specifically, when the reference model in DPO is set to a guiding reference model $\pi_{\text{guide}}$, the weight $\lambda$ becomes:
@@
The foresight of the guiding reference model is embodied in its ability to dynamically modulate $\lambda$, assigning higher weights to samples that the policy model is capable of learning well and down-weighting those that are difficult to learn or even conflict with the model’s learning trajectory. This behavior naturally satisfies the conditions suggested in the concurrent work (Gao et al., 2025), where avoiding overly difficult examples benefits alignment. Compared to the reference model used in standard DPO, which applies a fixed constraint with no foresight, the guiding reference model offers adaptive and data-dependent reweighting, leading to more efficient and targeted policy improvement.
@@ 5 Experiments
We empirically evaluate the effectiveness of Pre-DPO in enhancing existing preference optimization methods through a guiding reference model. To ensure a comprehensive and fair assessment, we conduct experiments on the Llama3.2 (Grattafiori et al., 2024) and Qwen2.5 (Yang et al., 2024) model series, including both Base and Instruct versions. We evaluate performance on two widely-used preference optimization benchmarks: AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v1.0 (Li et al., 2024). Given the sensitivity of preference optimization to hyperparameters (Meng et al., 2024), we follow SimPO’s experimental methodology and conduct an extensive hyperparameter search to ensure reliable results. All experiments are conducted based on the LlamaFactory (Zheng et al., 2024) repository, and all models and datasets used are publicly available.
@@ 5.1 Experimental setup
Models.
In our experiments, we primarily consider two widely recognized series of open-source models, Llama3.2-3B and Qwen2.5-7B, including both Base and Instruct versions. The diversity in model types and scales enables a more comprehensive evaluation of the method’s effectiveness.
Datasets.
For Base models, we obtain their corresponding SFT versions by training on the UltraChat-200k (Ding et al., 2023) dataset [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. Subsequently, the binarized version of the UltraFeedback (Cui et al., 2024) dataset [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized is utilized for preference optimization.
For Instruct models, we use them directly as SFT models. In the preference optimization stage, we construct two more on-policy preference datasets for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct, respectively. Specifically, we use prompts from the UltraFeedback dataset to sample responses from the model, generating 6 responses per prompt with a temperature of 0.8 and top-p of 0.95. Subsequently, following previous work(Meng et al., 2024), we utilize the ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024) reward model to score each response and select the highest-scoring and lowest-scoring responses to form preference pairs $(x,y^{+},y^{-})$.
@@
We obtain the guiding reference models for Pre-DPO based on two popular preference optimization methods: DPO (Rafailov et al., 2023), representing reference-based methods, and SimPO (Meng et al., 2024), representing reference-free methods. All preference optimization experiments use a batch size of 128, a maximum sequence length of 4096, and a cosine learning rate schedule with a 6% warmup ratio, training for 1 epoch. Given the importance of hyperparameter tuning in offline preference optimization (Meng et al., 2024), we perform extensive hyperparameter searches for all preference optimization experiments to ensure fairness. Specifically, for the key hyperparameters, including the learning rate, $\beta$ (for DPO and SimPO), and $\gamma$ (for SimPO), a two-stage tuning strategy is employed. We first fix the learning rate and search for the optimal $\beta$ or $\gamma$. Then, with the best $\beta$ or $\gamma$ fixed, we search for the optimal learning rate. More details of hyperparameter tuning and the best hyperparameter setting in our experiments can be found in Appendix A.1.
@@
We evaluate methods primarily on two open-source instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v0.1 (Li et al., 2024), which are widely adopted in the community for evaluating the instruction-following capabilities of LLMs (Meng et al., 2024). We report the raw win rate (WR) and the length-controlled win rate (LC) on AlpacaEval 2, and the win rate (WR) on Arena-Hard v0.1, with GPT-4-Turbo serving as the judge model. More evaluation details can be found in Appendix A.2.
@@ 5.2 Main results
Although Pre-DPO continuously improves performance, we observe that it does not significantly increase the response length compared to the baselines. Notably, in the Qwen2.5-7B-Base setting with SimPO as the guiding reference model, Pre-DPO even achieves the best performance with the shortest response length.
Additional observations.
Interestingly, we observe that SimPO shows a significant advantage over DPO in the Base settings, while DPO slightly outperforms SimPO in the Instruct settings. A possible reason is that the preference datasets constructed via on-policy sampling exhibit smaller length differences between positive and negative responses (see Appendix B.1 for details), thus diminishing the advantage of SimPO’s length normalization.
5.3 Ablations
@@
Policy Model Method Ref. Epochs AlpacaEval 2
LC (%) WR (%) Len.
Llama3.2-3B-Base-SFT DPO SFT 1 10.5 12.0 2042
DPO SFT 2 11.0 12.0 1976
Pre-DPO $\text{DPO}_{1}$ 1 12.5 13.9 2061
Llama3.2-3B-Instruct DPO SFT 1 36.3 36.9 2026
DPO SFT 2 35.2 37.1 2113
Pre-DPO $\text{DPO}_{1}$ 1 39.3 40.9 2095
Table 2: Ablation studies of the guiding reference model. DPO trained for 2 epochs has the same computational cost as Pre-DPO. $\text{DPO}_{1}$ denote the guiding reference model trained with DPO for 1 epoch.
@@
Although Pre-DPO consistently improves the performance of DPO and SimPO, it introduces additional computational cost due to the need to obtain a guiding reference model. To investigate whether the performance gain is simply due to increased training, we compare it with a baseline that trains DPO for 2 epochs with the original reference setting. As shown in Table 2, under the same computational budget, Pre-DPO with a guiding reference model consistently achieves the best LC and WR on AlpacaEval 2, benefiting from the better data utilization enabled by the guiding reference model and avoiding the excessive constraints imposed by traditional reference model setups.
Guiding reference models perform better on previously encountered preference data.
Method Ref. Training Set AlpacaEval 2
LC (%) WR (%) Len.
DPO SFT A 11.6 10.5 1702
Pre-DPO $\text{DPO}_{\text{A}}$ A 13.9 13.2 1819
Pre-DPO $\text{DPO}_{\text{B}}$ A 12.9 12.5 1791
DPO SFT B 13.9 12.9 1765
Pre-DPO $\text{DPO}_{\text{B}}$ B 14.1 13.4 1792
Pre-DPO $\text{DPO}_{\text{A}}$ B 13.8 12.7 1708
Table 3: Ablation studies of Pre-DPO under different guiding reference models and preference datasets. The SFT model is obtained by training Qwen2.5-3B-Base on the UltraChat-200k dataset. The preference training sets A and B are created by randomly splitting the UltraFeedback dataset into two equal parts. $\text{DPO}_{\text{A}}$ and $\text{DPO}_{\text{B}}$ denote the guiding reference models trained on training sets A and B, respectively.
To investigate the optimal relationship between the guiding reference model and the preference training data, we divide the preference dataset into two equal parts, A and B. We then perform DPO on each part separately to obtain two distinct guiding reference models. For each subset of preference data, we apply Pre-DPO using both guiding reference models for comparison. The results, as shown in Table 3, indicate that a guiding reference model trained on the same subset of preference data yields better performance on AlpacaEval 2. This suggests that guiding reference models perform better when guiding policies on data they have previously encountered, as they possess more accurate information about potential policy optimization directions and enable more effective data weighting.
5.4 The $\lambda$ distribution after training of Pre-DPO and vanilla DPO
(a) Llama3.2-3B-Base
(b) Llama3.2-3B-Instruct
Figure 3: Comparative differences in $\lambda$ (when $\beta$ is set to 0.01) distributions between Pre-DPO and DPO across Llama3.2-3B settings. Each subplot presents a violin plot illustrating the distribution of $\lambda$ values on the training preference dataset. The black horizontal line indicates the mean, while the dark gray line represents the median. The width of the violin at each point represents the data density, where thicker sections correspond to regions with higher data concentration.
Based on Eq. 5, Pre-DPO demonstrates a distinct advantage through its adaptive data reweighting mechanism, which dynamically assigns higher weights to preference data that align with the learned outcomes of the guiding reference model while diminishing the influence (assign lower weights) of data that may introduce contradictory learning signals. Figure 3 illustrates the differences in $\lambda$ distributions on the training preference dataset between Pre-DPO and DPO on Llama3.2-3B settings. For clarity, we report the statistics of the static $\lambda$ values computed on the training dataset after training completion. Pre-DPO produces higher mean and median $\lambda$ values compared to DPO, indicating that the guiding reference model effectively guides the policy model toward better data utilization. More comparisons between Pre-DPO and the baselines are provided in Appendix B.2.
@@ 6 Conclusion
We proposed Pre-DPO, a simple yet effective DPO-based preference optimization paradigm that enhances data utilization and improves performance by leveraging a guiding reference model. Unlike traditional DPO, which uses a reference identical to the initial policy, Pre-DPO reuses a previously optimized guiding reference model to re-optimize the initial policy model. This shifts the reference model’s role from constraint to guidance, enabling better data reweighting. Extensive experiments demonstrate that Pre-DPO consistently outperforms both DPO and SimPO across various models and scales, without relying on external models or additional data. We hope this work can inspire more exploration and discussion on the role and improvement of reference models in LLM alignment.
Appendix A More experimental details
A.1 Training hyperparameters
@@
Considering that hyperparameter tuning is crucial for preference optimization, we adopt a two-stage hyperparameter search method to ensure both fairness and efficiency. Specifically, in the first stage, we fix the learning rate at $6\times 10^{-7}$ and individually search for the optimal $\beta$ in $[0.005,0.01,0.05,0.1,0.2,0.5,1.0]$ for DPO, and search for $\beta$ in $[2.0,2.5]$ and $\gamma$ in $[0.3,0.5,1.0,1.2,1.4,1.6]$ for SimPO. In the second stage, we select the two best hyperparameter settings found in the first stage and individually search for the optimal learning rate in $[3\times 10^{-7},5\times 10^{-7},8\times 10^{-7},1\times 10^{-6}]$. The search ranges for the hyperparameters are chosen with reference to prior work (Meng et al., 2024).
@@
Table 4: Optimal hyperparameters for different experiments.
A.2 Evaluation details
Following previous work (Meng et al., 2024), we adopt a sampling-based decoding strategy for AlpacaEval 2 with a temperature of 0.7, a top-p of 0.9, and a greedy decoding strategy for Arena-Hard v0.1. We evaluate the AlpacaEval 2 and Arena-Hard v0.1 results using their official repositories, respectively.
Appendix B More analysis
B.1 Response length differences

### v3

@@ 4.2 Pre-DPO: Improving Data Utilization in DPO Using a Guiding Reference Model
Notably, a model that has already undergone preference optimization contains information about the entire training process. More importantly, it reflects the outcomes that the initial policy can achieve through the available preference data. Specifically, when the reference model in DPO is set to a guiding reference model $\pi_{\text{guide}}$, the weight $\lambda$ becomes:
@@
The foresight of the guiding reference model is reflected in its way to modulate $\lambda$: assigning higher weights to samples that the policy model can learn effectively, while down-weighting those that are difficult to learn or potentially conflicting. This behavior naturally aligns with findings suggesting that avoiding ambiguous, mislabeled, or overly difficult preference data can benefit alignment (houliston2024uncertainty; gao2025principled). Compared to the reference model used in standard DPO, which merely serves as a constraint without foresight, the guiding reference model enables more informed and data-dependent reweighting, leading to more efficient and targeted policy improvement.
@@ 5 Experiments
We empirically evaluate the effectiveness of Pre-DPO in enhancing existing preference optimization methods through a guiding reference model. To ensure a comprehensive and fair assessment, we conduct experiments on the Llama3.2-3B (grattafiori2024llama) and Qwen2.5-7B (yang2024qwen2) model series, including both Base and Instruct versions. We evaluate the models’ performance on two widely-used preference optimization benchmarks: AlpacaEval 2 (alpaca_eval; dubois2024length) and Arena-Hard v1.0 (li2024crowdsourced). Given the sensitivity of preference optimization to hyperparameters (meng2024simpo), we conduct an extensive hyperparameter search to ensure reliable results. All experiments are conducted based on the LlamaFactory (zheng2024llamafactory) repository, and all models and datasets used are publicly available.
@@ 5.1 Experimental Setup
Models and datasets.
In our experiments, we primarily consider two widely recognized series of open-source models, Llama3.2-3B and Qwen2.5-7B, including both Base and Instruct versions. The diversity in model types and scales enables a more comprehensive evaluation of our method’s effectiveness.
For Base models, we first train them on the UltraChat-200k (ding2023enhancing) dataset to obtain their corresponding SFT models. We then directly perform preference optimization on the existing binarized UltraFeedback (cui2024ultrafeedbackboostinglanguagemodels) dataset, which is widely used in prior work on offline preference optimization training (ethayarajh2024kto; hong2024orpo; meng2024simpo; liu2024understanding; kim2025sdpo; gorbatovski2025learn), using the SFT model as the initialization.
For Instruct models, we directly use them as SFT models. During the preference optimization stage, we construct two additional on-policy preference datasets for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct, respectively. Specifically, for each prompt in the binarized UltraFeedback dataset, we sample six responses from each Instruct model using a temperature of 0.8 and a top-$p$ of 0.95 as sampling parameters. Subsequently, following prior work(meng2024simpo), we use the ArmoRM-Llama3-8B-v0.1 (wang2024interpretable) reward model to score each response and select the highest-scoring and lowest-scoring responses to form preference pairs $(x,y^{+},y^{-})$. We discard prompts for which all sampled responses receive identical scores. This process results in two new preference datasets, one for each of the Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct models.
Baselines.
Prior work (meng2024simpo) shows that, with sufficient hyperparameter tuning, both DPO (rafailov2023direct) and SimPO (meng2024simpo) are highly competitive. Therefore, we adopt them as representative baselines for the reference-based and reference-free methods in our large-scale main experiments. Additionally, we conduct experiments with ORPO (hong2024orpo), sDPO (kim2025sdpo) and TR-DPO (gorbatovski2025learn) under the Llama3.2-3B-Base setting.
@@
All preference optimization experiments use a batch size of 128, a maximum sequence length of 4096, and a cosine learning rate schedule with a 6% warmup ratio, training for 1 epoch. All models are fine-tuned using full parameter tuning.
Given the importance of hyperparameter tuning in offline preference optimization (meng2024simpo), we perform extensive hyperparameter searches for all preference optimization experiments to ensure fairness. Specifically, for the key hyperparameters, including the learning rate, $\beta$ (for DPO-based methods and SimPO), and $\gamma$ (for SimPO), a two-stage tuning strategy is employed. We first fix the learning rate and search for the optimal $\beta$ or $\gamma$. Then, with the best $\beta$ or $\gamma$ fixed, we search for the optimal learning rate. More details of hyperparameter tuning and the best hyperparameter setting can be found in Appendix B.1.
@@
We evaluate methods primarily on two open-source instruction-following benchmarks: AlpacaEval 2 (alpaca_eval; dubois2024length) and Arena-Hard v0.1 (li2024crowdsourced), which are widely adopted in the community for evaluating the instruction-following capabilities of LLMs (meng2024simpo). We report the raw win rate (WR) and length-controlled win rate (LC) on AlpacaEval 2, and the WR on Arena-Hard v0.1, using their respective official repositories. More evaluation details can be found in Appendix B.2.
@@ 5.2 Main Results
Method Ref. Epoch AlpacaEval 2
LC (%) WR (%) Len.
Base DPO SFT 1 10.5 12.0 2042
DPO SFT 2 11.0 (+4.8%) 12.0 (+0.0%) 1976
Pre-DPO $\text{DPO}_{1}$ 1 12.5 (+19.0%) 13.9 (+15.8%) 2061
Instruct DPO SFT 1 36.3 36.9 2026
DPO SFT 2 35.2 (-3.0%) 37.1 (+0.5%) 2113
Pre-DPO $\text{DPO}_{1}$ 1 39.3 (+8.3%) 40.9 (+10.8%) 2095
Table 2: Ablation studies under the Llama3.2-3B model settings. DPO trained for 2 epochs has the same computational cost as Pre-DPO. $\text{DPO}_{1}$ denote the guiding reference model trained with DPO for 1 epoch.
Method Ref. AlpacaEval 2
LC (%) WR (%) Len.
ORPO - 10.2 7.9 1588
Pre-DPO ORPO 12.3 (+20.6%) 12.1 (+53.2%) 1907
sDPO last stage 12.0 11.9 1908
Pre-DPO sDPO 12.9 (+7.5%) 13.0 (+9.2%) 1951
TR-DPO hard update 11.7 12.3 1985
Pre-DPO TR-DPO 12.8 (+9.4%) 14.2 (+15.4%) 2087
Table 3: More results of Pre-DPO with diverse guiding reference models under Llama3.2-3B-Base setting.
@@
Although Pre-DPO continuously improves performance, we observe that it does not significantly increase the response length compared to the baselines. Notably, with SimPO as the guiding reference model, Pre-DPO achieves the best performance and the shortest average response length in the Qwen2.5-7B-Base setting.
Pre-DPO is compatible with the iterative preference optimization framework.
Note that the DPO and SimPO experiments for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct use on-policy preference datasets constructed by sampling from the current policy, corresponding to the first round of iterative preference optimization (xiong2024iterative; yuan2024self; rosset2024direct; zhang2025iterative). Under this setting, Pre-DPO achieves better optimization performance, indicating that Pre-DPO is complementary to the iterative framework and can be employed as part of the iterative process to enhance the use of newly collected preference data.
5.3 Ablations and More Results
@@
Figure 3: $\lambda$ distribution dynamics of DPO, TR-DPO, and Pre-DPO under the Llama3.2-3B-Base setting. Pre-DPO maintains a broader distribution during the entire training.
Figure 4: Quantitative analysis of the $\lambda$ distribution during training for DPO, TR-DPO (hard update), and Pre-DPO under the Llama3.2-3B-Base setting. Numerical values on top of the bars indicate the corresponding percentages.
@@
Although Pre-DPO consistently improves the performance of DPO and SimPO, it introduces additional computational cost due to the need to obtain a guiding reference model. To investigate whether the performance gain is simply due to increased training, we compare Pre-DPO with a baseline where DPO is trained for 2 epochs using the original reference configuration. As shown in Table 2, DPO with a larger computational budget does not yield a noticeable gain, which aligns with previous findings that a single training epoch generally yields the best results (meng2024simpo). However, Pre-DPO with a guiding reference model consistently achieves the best LC and WR on AlpacaEval 2, benefiting from the better data utilization enabled by the guiding reference model and avoiding the excessive constraints imposed by traditional reference model setups.
Pre-DPO can consistently benefit from diverse preference optimization methods.
To further validate the generality of Pre-DPO, we obtain guiding reference models from more preference optimization methods under the Llama3.2-3B-Base setting, including ORPO, sDPO, and TR-DPO. For sDPO, we divide the preference dataset into two equal parts and perform a two-stage DPO training procedure. For TR-DPO, we adopt a hard update strategy, where the reference model is reset to the current policy every 32 training steps. As shown in Table 3, Pre-DPO consistently improves performance on AlpacaEval 2. The results demonstrate that by leveraging guiding reference models, Pre-DPO can consistently benefit from diverse preference optimization methods.
5.4 The $\lambda$ Distribution during Pre-DPO Training
Pre-DPO demonstrates a clear advantage via adaptive data reweighting, assigning higher weights to samples aligned with the guiding reference model and down-weighting those that may introduce conflicting learning signals.
To better illustrate the training dynamics, we visualize the evolution of the $\lambda$ distribution throughout training for DPO, TR-DPO, and Pre-DPO under the Llama3.2-3B-Base setting. As intuitively observed in Figure 3, Pre-DPO, equipped with a guiding reference model, maintains a broader $\lambda$ distribution throughout training. In contrast, DPO initially assigns nearly uniform data weights and gradually adjusts $\lambda$ to regulate the policy’s deviation from the initial policy. TR-DPO with hard reference updates (i.e., resetting the reference model to the current policy every $k$ steps) tends to maintain a more uniform $\lambda$ distribution throughout training, thereby diminishing the influence of the reference model.
Figure 4 quantitatively demonstrates that in the early stages of training (0–33%), Pre-DPO’s $\lambda$ values are more concentrated at the extremes (i.e., in the ranges 0–0.2 and 0.8–1.0), whereas those of DPO and TR-DPO are more centered around 0.5. As training progresses into the later stages (34–100%), Pre-DPO’s $\lambda$ distribution shifts closer to 0.5. Nevertheless, it consistently maintains a broader and more balanced distribution across the entire training.
@@ 6 Conclusion
We proposed Pre-DPO, a simple yet effective DPO-based preference optimization paradigm that enhances data utilization and improves performance by leveraging a guiding reference model. Unlike traditional DPO, which uses a reference identical to the initial policy, Pre-DPO reuses an optimized policy model as the guiding reference model to re-optimize the initial policy model. This shifts the role of the reference model from a constraint to an informed guide, enabling more effective data reweighting. Extensive experiments across multiple models and scales show that Pre-DPO consistently outperforms both DPO and SimPO, without requiring external models or additional data. We hope this work can inspire more exploration and discussion on the role and improvement of reference models in RLHF for LLMs.
Acknowledgments
We would like to sincerely thank all the anonymous reviewers for their valuable feedback. This work is partially funded by the National Natural Science Foundation of China Key Program (Grant No. 62336006).
Appendix A Derivation of the DPO Gradient
Given the DPO loss
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})\!=\!$ $-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\Big(\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}$
$\qquad\qquad-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\Big)\Big],$ (11)
let
$z=\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\text{ref}}(y^{+}|x)}-\beta\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)},$ (12)
the gradient can be expressed as
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\frac{1}{\sigma(z)}\sigma^{\prime}(z)\nabla_{\theta}z\right]$
$=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[(1-\sigma(z))\nabla_{\theta}z\right]$
$=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\sigma(-z)\nabla_{\theta}z\right]$
$=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\sigma\Big(\beta\log\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}$
$-\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\theta}(y^{-}|x)}\Big)\nabla_{\theta}\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\theta}(y^{-}\mid x)}\Big].$ (13)
We define:
$\lambda=\sigma\left(\beta\log\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}-\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\theta}(y^{-}|x)}\right),$ (14)
and thus we can obtain
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\beta\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\lambda\cdot\nabla_{\theta}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\theta}(y^{-}\mid x)}\right].$ (15)
Appendix B More Experimental Details
B.1 Training Hyperparameters
@@
Considering that hyperparameter tuning is crucial for preference optimization, we adopt a two-stage hyperparameter search method to ensure both fairness and efficiency. Specifically, in the first stage, we fix the learning rate at $6\times 10^{-7}$ and individually search for the optimal $\beta$ in $[0.005,0.01,0.05,0.1,0.2,0.5,1.0]$ for DPO-based methods, search for $\beta$ in $[2.0,2.5]$ and $\gamma$ in $[0.3,0.5,1.0,1.2,1.4,1.6]$ for SimPO, and search for $\lambda$ in $[0.1,0.5,1.0,2.0,3.0,5.0]$ for ORPO. In the second stage, we select the two best hyperparameter settings found in the first stage and individually search for the optimal learning rate in $3\times 10^{-7},5\times 10^{-7},8\times 10^{-7},1\times 10^{-6}]$. The search ranges for the hyperparameters are chosen with reference to prior work (meng2024simpo).
@@
Table 4: Optimal hyperparameters for the main experiments.
B.2 Evaluation Details
Following previous work (meng2024simpo), we adopt a sampling-based decoding strategy for AlpacaEval 2 with a temperature of 0.7, a top-$p$ of 0.9, and a greedy decoding strategy for Arena-Hard v0.1. We evaluate the AlpacaEval 2 and Arena-Hard v0.1 results using their official repositories, both of which employ gpt-4-1106-preview as the judge model.
Appendix C More Analysis
C.1 Response Length Differences of $y^{+/-}$ Across Datasets
It can be observed that SimPO shows a significant advantage over DPO in the Base settings, while DPO slightly outperforms SimPO in the Instruct settings. A possible reason is that the preference datasets constructed via on-policy sampling exhibit smaller length differences between positive responses $y^{+}$ and negative responses $y^{-}$, thus diminishing the advantage of SimPO’s length normalization.